OpenStudy (anonymous):

Find S1, S2, S3, S4, S5 and use these sums to approximate S. Then use the formula to find S. 8 + 2 + 1/2 + 1/8 + ...

OpenStudy (dumbcow):

looks like a geometric sequence with r = 1/4

OpenStudy (dumbcow):

formula for infinite geometric sequence is: \[S = \frac{a}{1-r} = \frac{8}{1-\frac{1}{4}}\]

OpenStudy (anonymous):

how do you find the S4 and S5?

OpenStudy (dumbcow):

sorry i don't know what S1,S2... refer to